OpenStudy (anonymous):

Solve the Equation 4d+2d-4d+4=3d A.) 4/7 B.) 4 C.) -4/5 D.) -4

OpenStudy (ryan123345):

combine all the like terms first. What do u get?

OpenStudy (anonymous):

10d

OpenStudy (anonymous):

nope! try again. 4d-4d=0

OpenStudy (ryan123345):

what is 4d plus 2d?

OpenStudy (anonymous):

This equation is fairly simple. you can either add 4d+2d=6d 6d-4d=2d or simply cancel out 4d and -4d which leaves you with 2d.

OpenStudy (anonymous):

what part do you not understand?

OpenStudy (anonymous):

Keep in mind you are solving for D. so you want to get d= to something.